# UFO [Uncaptioned image]: A UI-Focused Agent for Windows OS Interaction

### v4

@@ 6 Conclusion
We introduce UFO [Uncaptioned image], an innovative UI-focused agent designed to fulfill user requests in natural language through intelligent interactions with applications on the Windows OS. Leveraging GPT-Vision, UFO analyzes application GUI screenshots and controls information to dynamically select the optimal application and control for executing actions and satisfying user queries. The grounding of actions on the application is facilitated by a control interaction module, minimizing human intervention and enabling full automation, establishing it as a comprehensive LAM framework. With the dual-agent design, namely, AppAgent and ActAgent, UFO intelligently switches between applications, allowing for the completion of long and complex tasks that span different applications. Additionally, we incorporate key features such as Action Customization and Safeguard to enhance UFO’s extendibility and safety. Evaluation results based on 50 requests across 9 popular Windows OS applications demonstrate its remarkable versatility and generalizability. To the best of our knowledge, UFO represents the pioneering effort in developing a UI automation agent specifically tailored for the Windows OS environment.
@@ Appendix A Requests in WindoesBench and Detailed Evaluations
I am Zac. Draft an email, to [email address 1] and cc [email address 2], to thanks for his contribution on the VLDB paper on using diffusion models for time series anomaly detection. Don’t send it out.
Outlook
6
100%
Search ’Spring Festival’ in my mail box, and open the second returned email.
Outlook
4
100%
Delete the first junk email.
Outlook 4
100%
Please download and save the pdf attachment to the local in the first sent email.
Outlook
10
70%
(1) Draft an email, to {email address 1} and to ask him if he can come to the meeting at 3:00 pm. Don’t send it out.
(2) Add some content to the email body to tell him the meeting is very important, and cc {email address 2} as well.
(3) Send the email now.
Outlook 10
100%
Give a detailed description of the components and workflows of the TaskWeaver architecture in the image.
Photos
2
100%
Open the image of the LLM training and rotate it for 2 times.
Photos
3
100%
Open the image of the LLM training and delete it.
Photos 4
100%
Open the mouse image, then loop over the next two images, without closing them. Summarize them one by one.
Photos
5
83.3%
(1) Zoom in the image of the autogen for 2 times to make it clearer, then describe it with detailed workflow and components.
(2) It looks too big. Just zoom it to fit the screen.
Photos
6
100%
Help me quickly remove all notes in the slide of the ufo_testing, without looping through each slide one-by-one.
PowerPoint 8
100%
Clear the recording on current page for ppt.
PowerPoint 3
100%
Please add Morph transition to the current page and the next page of the ufo_testing.ppt
PowerPoint
6
83.3%
Summarize scripts for all pages in the ppt one by one to help with my presetation.
PowerPoint
7
85.7%
(1) Apply the first format generated by the Designer to the current ppt.
(2) This format is not beautiful. Select a better format and explain why you select it.
PowerPoint
8
75%
Check spelling and grammar of the current meeting note.
Word
2
100%
Please change the theme of the meeting note into ’Organic’.
Word
4
100%
Save meeting note as an Adobe PDF to the local.
Word
6
83.3%
Add a cover page to the meeting note. Choose one you think is beautiful.
Word
5
100%
(1) Please change to a better-looking format and color to beautify the current page of the meeting note.
(2) Do more to make it look even better.
Word
10
80%
@@
Close all pdf that are currently open.
Adobe Acrobat 3
100%
Cascade the current pdf windows and close the first one.
Adobe Acrobat 3
60%
Visually find and present to me the timeline and milestones in the current meeting note.
Adobe Acrobat
4
50%
(1) Read the imdiffuion paper. What is the main contribution of it?
(2) Find the first figure of the paper. What does it show?
Adobe Acrobat
5
100%
(1) Comprehend the current pdf, what is the difference between phase 1 and phase 2?
(2) What is the difference of the Application Selection Agent and the Action Selection Agent?
(3) Base on the workflow pdf, is there anything can be improved for the agents?
(4) Thanks, this is actually the workflow of yourself. Do you think there are other things can be improved for you? I will implement your thought on it.
(5) Great. Do you think there are other things that can be improved visually for the figure itself in the pdf, to provide better illustration for other readers?
Adobe Acrobat
6
83.3%
Navigate to /Desktop/mix/screen and open s0.png by only double clicking.
File Explorer
6
83.3%
Delete all files individually in the current figure folder.
File Explorer 6
100%
Copy the actagent file in the folder to the Document folder.
File Explorer
5
60%
Create a new txt file in the current folder.
File Explorer
3
100%
(1) Open all images that are related to dragon in the test folder.
(2) Summarize in detailed what you have seen in the image d4.
File Explorer
4
100%
Tell me the changelog of the {code base name} repo in their readme base on your observation.
Visual Studio Code
5
80%
Find the use of ’print_with_color’ in the script folder in the {code base name} repo.
Visual Studio Code
4
100%
Download the Docker extension in the {code base name} repo.
Visual Studio Code
4
100%
Create a new terminal at the {code base name} repo.
Visual Studio Code
2
100%
(1) Read through carefully and review the code in the utils.py in the {code base name} and identify any potential bugs or aspects can be improved.
(2) What about in the draw_bbox_multi function?
Visual Studio Code
5
40%
Please send a ’smile’ emoticon to the current chatbox at Wechat.
WeChat 4
100%
Delete the current chatbox of ’File Transfer’ on Wechat.
WeChat 3
100%
Please like {user name}’s post at Moment.
WeChat
3
100%
Bring the first picture to the front in my ”favourite” at WeChat and decribe it.
WeChat
5
100%
(1) Open the {account name} Official Account at WeChat.
(2) Send a message to ’File Transfer’ to introduce this account.
WeChat 10
90%
@@
How many total citations does Geoffrey Hinton have currently?
Edge Browser
4
100%
Post ”It’s a good day.” on my Twitter.
Edge Browser
6
100%
Download the Imdiffusion repo as zip.
Edge Browser
7
100%
Change the theme of the browser to the icy mint.
Edge Browser
5
100%
(1) Find and navigate to {name}’s homepage at Microsoft.
(2) Print this page in color.
Edge Browser
5
60%
My name is Zac. Please read the meeting note in meeting_note.docx to identify all action items, and include the detailed description LLM training workflow in the LLM-training.png, to compose an new email of these contents. Sent the full email to our leader Hidan via {email address} to ask for his review.
Word, Photos, Outlook
10
100%
Search for and read through the latest news about Microsoft and summarize it, then send it to ’File Transfer’ on WeChat.
Edge Browser, WeChat
9
100%
Open the image of ufo_rv at UFO-windows/logo/, summarize its content and use the summary to search for a similar image on Google.
File Explorer, Edge Browser
10
90%
Observe the overview.pdf, then download the most similar paper base on this architecture from the Internet base on your understanding, and sent it to the ’File Transfer’ on WeChat.
Adobe Acrobat, Edge Browser, WeChat
10
25%
(1) Read the paper.pptx, summarize the paper title presented on the slide, then search and open the paper on the Internet.
(2) Can you summarize this paper and download its PDF version for me?
PowerPoint, Edge Broswer
10
100%
@@
Dear Hidan,
Please find below the action items identified from our recent meeting, as well as a detailed description of the LLM training workflow.
Action Items:
- Zac Yang: Develop a comprehensive project plan and share it with the team by 2024.03.06. - Kevin Cheung: Conduct thorough research on potential risks associated with LLM training and present findings by 2024.04.01. - Tom James: Formulate a detailed proposal for resource allocation, including computing resources and software tools. Submit it for review by 2024.01.03. - Sophia Chow: Compile and distribute meeting minutes, including action items and deadlines, within 24 hours.
LLM Training Workflow Description: The LLM training framework consists of four main components: ’Pre-training’ with no human involvement using proprietary and unlabeled data, ’Self-supervised learning’ where the base model is trained with narrow datasets, ’Fine-tuning’ which involves human feedback to fine-tune the model, and ’Prompt-based learning’ where the augmented model is developed through prompting with human involvement and specialized knowledge.
Please review the above information and provide your feedback at your earliest convenience.
Best regards, Zac

### v5

@@ 6 Conclusion
We introduce UFO [Uncaptioned image], an innovative UI-focused agent designed to fulfill user requests in natural language through intelligent interactions with applications on the Windows OS. Leveraging GPT-Vision, UFO analyzes application GUI screenshots and controls information to dynamically select the optimal application and control for executing actions and satisfying user queries. The grounding of actions on the application is facilitated by a control interaction module, minimizing human intervention and enabling full automation, establishing it as a comprehensive LAM framework. With the dual-agent design, namely, HostAgent and AppAgent, UFO intelligently switches between applications, allowing for the completion of long and complex tasks that span different applications. Additionally, we incorporate key features such as Action Customization and Safeguard to enhance UFO’s extendibility and safety. Evaluation results based on 50 requests across 9 popular Windows OS applications demonstrate its remarkable versatility and generalizability. To the best of our knowledge, UFO represents the pioneering effort in developing a UI automation agent specifically tailored for the Windows OS environment.
@@ Appendix A Requests in WindoesBench and Detailed Evaluations
I am Zac. Draft an email, to [email address 1] and cc [email address 2], to thanks for his contribution on the VLDB paper on using diffusion models for time series anomaly detection. Don’t send it out. Outlook - 6 100%
Search ’Spring Festival’ in my mail box, and open the second returned email. Outlook - 4 100%
Delete the first junk email. Outlook 4 100%
Please download and save the pdf attachment to the local in the first sent email. Outlook - 10 70%
(1) Draft an email, to {email address 1} and to ask him if he can come to the meeting at 3:00 pm. Don’t send it out. (2) Add some content to the email body to tell him the meeting is very important, and cc {email address 2} as well. (3) Send the email now. Outlook 10 100%
Give a detailed description of the components and workflows of the TaskWeaver architecture in the image. Photos - 2 100%
Open the image of the LLM training and rotate it for 2 times. Photos - 3 100%
Open the image of the LLM training and delete it. Photos 4 100%
Open the mouse image, then loop over the next two images, without closing them. Summarize them one by one. Photos - 5 83.3%
(1) Zoom in the image of the autogen for 2 times to make it clearer, then describe it with detailed workflow and components. (2) It looks too big. Just zoom it to fit the screen. Photos - 6 100%
Help me quickly remove all notes in the slide of the ufo_testing, without looping through each slide one-by-one. PowerPoint 8 100%
Clear the recording on current page for ppt. PowerPoint 3 100%
Please add Morph transition to the current page and the next page of the ufo_testing.ppt PowerPoint - 6 83.3%
Summarize scripts for all pages in the ppt one by one to help with my presetation. PowerPoint - 7 85.7%
(1) Apply the first format generated by the Designer to the current ppt. (2) This format is not beautiful. Select a better format and explain why you select it. PowerPoint - 8 75%
Check spelling and grammar of the current meeting note. Word - 2 100%
Please change the theme of the meeting note into ’Organic’. Word - 4 100%
Save meeting note as an Adobe PDF to the local. Word - 6 83.3%
Add a cover page to the meeting note. Choose one you think is beautiful. Word - 5 100%
(1) Please change to a better-looking format and color to beautify the current page of the meeting note. (2) Do more to make it look even better. Word - 10 80%
@@
Close all pdf that are currently open. Adobe Acrobat 3 100%
Cascade the current pdf windows and close the first one. Adobe Acrobat 3 60%
Visually find and present to me the timeline and milestones in the current meeting note. Adobe Acrobat - 4 50%
(1) Read the imdiffuion paper. What is the main contribution of it? (2) Find the first figure of the paper. What does it show? Adobe Acrobat - 5 100%
(1) Comprehend the current pdf, what is the difference between phase 1 and phase 2? (2) What is the difference of the Application Selection Agent and the Action Selection Agent? (3) Base on the workflow pdf, is there anything can be improved for the agents? (4) Thanks, this is actually the workflow of yourself. Do you think there are other things can be improved for you? I will implement your thought on it. (5) Great. Do you think there are other things that can be improved visually for the figure itself in the pdf, to provide better illustration for other readers? Adobe Acrobat - 6 83.3%
Navigate to /Desktop/mix/screen and open s0.png by only double clicking. File Explorer - 6 83.3%
Delete all files individually in the current figure folder. File Explorer 6 100%
Copy the actagent file in the folder to the Document folder. File Explorer - 5 60%
Create a new txt file in the current folder. File Explorer - 3 100%
(1) Open all images that are related to dragon in the test folder. (2) Summarize in detailed what you have seen in the image d4. File Explorer - 4 100%
Tell me the changelog of the {code base name} repo in their readme base on your observation. Visual Studio Code - 5 80%
Find the use of ’print_with_color’ in the script folder in the {code base name} repo. Visual Studio Code - 4 100%
Download the Docker extension in the {code base name} repo. Visual Studio Code - 4 100%
Create a new terminal at the {code base name} repo. Visual Studio Code - 2 100%
(1) Read through carefully and review the code in the utils.py in the {code base name} and identify any potential bugs or aspects can be improved. (2) What about in the draw_bbox_multi function? Visual Studio Code - 5 40%
Please send a ’smile’ emoticon to the current chatbox at Wechat. WeChat 4 100%
Delete the current chatbox of ’File Transfer’ on Wechat. WeChat 3 100%
Please like {user name}’s post at Moment. WeChat - 3 100%
Bring the first picture to the front in my ”favourite” at WeChat and decribe it. WeChat - 5 100%
(1) Open the {account name} Official Account at WeChat. (2) Send a message to ’File Transfer’ to introduce this account. WeChat 10 90%
@@
How many total citations does Geoffrey Hinton have currently? Edge Browser - 4 100%
Post ”It’s a good day.” on my Twitter. Edge Browser 6 100%
Download the Imdiffusion repo as zip. Edge Browser - 7 100%
Change the theme of the browser to the icy mint. Edge Browser - 5 100%
(1) Find and navigate to {name}’s homepage at Microsoft. (2) Print this page in color. Edge Browser - 5 60%
My name is Zac. Please read the meeting note in meeting_note.docx to identify all action items, and include the detailed description LLM training workflow in the LLM-training.png, to compose an new email of these contents. Sent the full email to our leader Hidan via {email address} to ask for his review. Word, Photos, Outlook - 10 100%
Search for and read through the latest news about Microsoft and summarize it, then send it to ’File Transfer’ on WeChat. Edge Browser, WeChat 9 100%
Open the image of ufo_rv at UFO-windows/logo/, summarize its content and use the summary to search for a similar image on Google. File Explorer, Edge Browser - 10 90%
Observe the overview.pdf, then download the most similar paper base on this architecture from the Internet base on your understanding, and sent it to the ’File Transfer’ on WeChat. Adobe Acrobat, Edge Browser, WeChat - 10 25%
(1) Read the paper.pptx, summarize the paper title presented on the slide, then search and open the paper on the Internet. (2) Can you summarize this paper and download its PDF version for me? PowerPoint, Edge Broswer - 10 100%
@@
Dear Hidan, Please find below the action items identified from our recent meeting, as well as a detailed description of the LLM training workflow. Action Items: - Zac Yang: Develop a comprehensive project plan and share it with the team by 2024.03.06. - Kevin Cheung: Conduct thorough research on potential risks associated with LLM training and present findings by 2024.04.01. - Tom James: Formulate a detailed proposal for resource allocation, including computing resources and software tools. Submit it for review by 2024.01.03. - Sophia Chow: Compile and distribute meeting minutes, including action items and deadlines, within 24 hours. LLM Training Workflow Description: The LLM training framework consists of four main components: ’Pre-training’ with no human involvement using proprietary and unlabeled data, ’Self-supervised learning’ where the base model is trained with narrow datasets, ’Fine-tuning’ which involves human feedback to fine-tune the model, and ’Prompt-based learning’ where the augmented model is developed through prompting with human involvement and specialized knowledge. Please review the above information and provide your feedback at your earliest convenience. Best regards, Zac
@@ D.5 Sending the News
The Microsoft News page displays several articles. The headlines include: ’Excited to announce the general availability of Copilot for Sales and Copilot for Service, as we continue to extend Copilot to every role and function.’, ’Microsoft announces quarterly earnings release date’, ’With Copilot Pro, we’re helping even more people supercharge their creativity and productivity by unlocking Copilot in Microsoft 365 apps, providing access to the very latest models 2014 and more.’, and ’Microsoft unveils new generative AI and data solutions across the shopper journey, offering copilot experiences through Microsoft Cloud for Retail’.
@@ D.6 Search Paper in a PowerPoint Slide and Summarize
The paper titled ‘Xpert: Empowering Incident Management with Query Recommendations via Large Language Models’ discusses the role of large-scale cloud systems in modern IT infrastructure and the impact of incidents occurring within these systems. The paper introduces Xpert, a machine learning framework that automates Kusto Query Language (KQL) recommendations, tailored to new incidents. By leveraging historical incident data and large language models, Xpert generates customized KQL queries, enabling a thorough evaluation of query quality from three comprehensive perspectives. The paper details the deployment of Xpert in the real production environment of a large-scale incident management system in Microsoft, demonstrating its effectiveness in supporting incident management.